OpenStudy (anonymous):

HELP PLEASE ! What is the simplified form of the expression sqrt16c^4 - sqrt c^2 + 3sqrt c^2 + sqrt 4c^2

OpenStudy (anonymous):

\[\bf \sqrt{16c^4}=\sqrt{16}\sqrt{c^4}=4c^2\]

OpenStudy (anonymous):

You can deal with the rest of the expression in a similar manner.

OpenStudy (anonymous):

@snewton29 Try and do the rest now. Shouldn't be hard.

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

@genius12